Here is a set of sample data.

28 86 45 23 97 63 7 9 36 19
85 22 24 74 100 26 40 60 16 54
39 49 79 53 30

Base on the data above, identify the 5 number summary

_________, __________, __________, __________, __________
Answer #1

5 number summary:

Minimum = 7

First quartile, Q1 = 0.25(n+1)th value = 6.5th value of sorted data = 23.5

Median = 0.5(n+1)th value = 13th value of sorted data = 40

Third quartile, Q3 = 0.75(n+1)th value = 19.5th value of sorted data = 68.5

Maximum = 100